OpenStudy (anonymous):

n the sum sn sigma k^3/n^4 k=1 can be interpreted as a riemann sum for which integral

OpenStudy (anonymous):

first factor out a 1/n fron i^3/n^4 to get (i/n)^3 * 1/n Now, we know that delta x = b-a/n, so 1=b-a and xbasei* = a+ideltax so xbasei* = a+ideltax = a+(i/n)^3 so a =0 because a=0, b=1 , ill draw the solution in an attachment

OpenStudy (anonymous):

 is the answer